OpenStudy (anonymous):

A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing 24% solutions with a 50% solution to get a 36% solution. How many gallons of each will he use?

OpenStudy (anonymous):

Let x be the volume of the 24% solution. 24* x + 50* (100-x) = 36 * 100 24x + 5000 -50x = 3600 5000 - 3600 = 50x - 24x 1400 = 26x x = 53.8 Use 53.8 gallons of 24% solution and 100-53.8= 46.2 gallons of the 50% solution.